Plzzz say me how to differentiate x(x-3)^2 , I'm really confused plz solve this and say me........

Grade:11

2 Answers

Arnab Saha
7 years ago
You can solve it by two methods:1) Break the polynomial and then differentiate each variable separately 2) Use the product rule of differentiation taking u=x and v=(x-3)^2
Nandana
7 years ago
d/dx(X*(X-3)^2
this is like d/dx(uv)
take  u=X & v=(X-3)^2 , then
         = X*2*(X-3)*1+(X-3)^2*1
        = 2X^2-6X+X^2+9-6X
        = 3X^2-12X+9
        =3*(X^2-4X+3)